all surfaces are smooth,pulleys and strings are ideal. Masses of each block is 1 kg. if aA, aB are the acceleration of block a and b respectively then the value of aA is?


1
Expert's answer
2021-12-03T12:30:22-0500
"a_A=\frac{m_a-m_b}{m_a+m_b}g\\a_A=\frac{1-1}{1+1}9.8=0\frac{m}{s^2}"